OpenStudy (anonymous):

This year, John played in 10 baseball games. In these games he had hit the ball 2, 3, 0, 1, 3, 2, 4, 0, 2, and 3 times. In the first 10 games he plays next year, John wants to increase his average (mean) hits per game by 0.5. What is the total number of hits John needs over the first 10 games next year to achieve his goal?

OpenStudy (anonymous):

\[(2+3+0+1+3+2+4+0+2+3 ) / 10 = 2 , \[2*10 = 20 , 2+.5 * 10 = 25\]\] 25 is the total number he needs to hit to increase his mean by .5 or to 2.5

OpenStudy (anonymous):

sorta getting it now..

OpenStudy (lilai3):

congrats on ur first question

OpenStudy (anonymous):

still sorta confused tho

OpenStudy (anonymous):

the mean of year 1 is 20 divided by total games, which is 2, 10 games, the mean is 2 hits so you add .5 to it, thats how much John wants to increase by, which is from 2 to 2.5 hits, then to find the number of hits we can do this backwards, since we have the mean , 2.5, we can instead of dividing, we multiply 2.5 by 10 to get 25 hits in 10 games. Let me know if this helped

OpenStudy (anonymous):

yes i understand now thank you